A DI has the following assets in its portfolio: $31 million in cash reserves with the Fed, $31 million in T-bills, and $35 million in mortgage loans. If it needs to dispose of its assets at short notice, it will receive only 99 percent of the fair market value of the T-bills and 91 percent of the fair market value of its mortgage loans. If the DI waits one month to liquidate these assets, it would receive the full fair market value for each security. Calculate the one-month liquidity index using the previous information. (Do not round intermediate calculations. Round your answer to 3 decimal places. (e.g., 32.161))
Answer #1

DI's Assets: $ 31 million in Cash, $ 31 million in T-Bills and $ 35 million in Mortgage Loans

Total Assets = 31 + 31 + 35 = $ 97 million

Expected Receipt: 99% of the fair market value of the T-Bills and 91% of the fair market value of the mortgage loans.

Therefore, 1-Month Liquidity Index = (31/97) x 1 + (31/97) x 0.99 + (35/97) x 0.91 = 0.96433 ~ 0.964